### v1

@@ 4 Summary and discussion
While our current analysis suggests that PTAs cannot yet distinguish between isocurvature and adiabatic fluctuations, we anticipate that future advancements in observational capabilities and data analysis techniques will gradually improve our ability to probe the primordial origins of the SGWB. The continued progress in constructing next-generation radio telescopes, such as the Square Kilometre Array [133], promise to provide unprecedented sensitivity and frequency coverage, enabling more precise measurements of the SGWB and its properties.
@@
ZCC is supported by the National Natural Science Foundation of China (Grant Nos. 12247176 and 12247112) and the China Postdoctoral Science Foundation Fellowship (Grant No. 2022M710429). LL is supported by the National Natural Science Foundation of China (Grant Nos. 12247112 and 12247176) and the China Postdoctoral Science Foundation Fellowship (Grant No. 2023M730300).

### v2

@@ 5 Summary and discussion
While our current analysis suggests that PTAs cannot yet distinguish between isocurvature and adiabatic fluctuations, we anticipate that future advancements in observational capabilities and data analysis techniques will gradually improve our ability to probe the primordial origins of the SGWB. The continued progress in constructing next-generation radio telescopes, such as the Square Kilometre Array [141], promise to provide unprecedented sensitivity and frequency coverage, enabling more precise measurements of the SGWB and its properties. It is crucial to recognize that our analysis is based on the assumption of either purely adiabatic or purely isocurvature perturbations. However, the total fluctuations in reality may comprise a combination of both types of perturbations. While incorporating a ratio between the contributions of adiabatic and isocurvature perturbations to the IGW energy-density spectra is beyond the scope of the current study, it presents a promising direction for future research. We leave the exploration of mixed scenarios for subsequent studies, as they may offer valuable insights into the nature of primordial perturbations.
@@
Zu-Cheng Chen is supported by the National Natural Science Foundation of China under Grant No. 12405056 and the innovative research group of Hunan Province under Grant No. 2024JJ1006. Lang Liu is supported by the National Natural Science Foundation of China Grant under Grant No. 12433001.